OpenStudy (anonymous):

There were 68 family members at the picnic. Most of them came in minibuses. If 8 people could it in each minibus, how many minibuses did they need? How many people had to drive separately?

hero (hero):

8 *8 = 64 68 - 64 = 4

hero (hero):

8 minibuses 4 individual drivers

OpenStudy (anonymous):

How did you get the 2 8's

hero (hero):

Think division and remainder 68/8 = ?

OpenStudy (anonymous):

Okay so that the answer is 8 remainder of 4

hero (hero):

Precisely

hero (hero):

Feel free to award medals for good answers

OpenStudy (anonymous):

Thank you! I got it now.